What volume would 20.0 moles of sulfur dioxide occupy at 75.3 °C with a pressure of 3.14 atm?
Answer #1

Given:
P = 3.14 atm
n = 20.0 mol
T = 75.3 oC
= (75.3+273) K
= 348.3 K

use:
P * V = n*R*T
3.14 atm * V = 20 mol* 0.08206 atm.L/mol.K * 348.3 K
V = 182 L
Answer: 182 L